Fermi National Laboratory


Accelerator Update

Wednesday May 26
The day shift began with store #3532 in the Tevatron and colliding. The Pbar stack was at 174 mA. Recycler was doing studies and MiniBooNe was taking beam.

At 10:03 AM MCenter was running and the tune was documented.

Since the stack of PBars was pretty high, stacking was stopped for about an hour of slipstacking studies.

At 10:30 CDF requested the operators look at the abort gap losses, they were tuned and reduced.

Tevatron end of store studies begins at 11:00 AM.

At 13:00 for about an hour, Tevatron alpha bumps were changed and the experiments took data with the new positions for an hour.

At 14:30 the Meson Test experiment was secure so they could begin taking beam.

At 14:45 the Tevatron store was terminated and beam to Meson Test followed.

The C3 wet engine needed some maintenance, so time was allowed between the stores.

At 16:20, shot setup began.

KRF3 tripped off while loading final protons, it was turned on and loading continued.

PBars were then loaded.

Once the manipulations were done, to begin store 3534, the initial luminosity came out to be 73.3 e30. This is a new record!

Some tuning was done to reduce losses and the experiments were taking data.

At 20:40, shot setup for PBars to the recycler began and was complete within an hour.

By 21:45, stacking continued.

MI-13 was secured after a MiniBooNe access and beam was restored within 10 minutes.

The machines continued to run well the remainder of the shift.

Thursday May 27
The midnight shift began with store 3534 in the TeV. PBar was stacking. MiniBooNe was taking beam. MCenter and MTest were also taking beam.

Things were running pretty well until about 05:33 AM when a power glitch occurred.

The Tevatron store was lost and all machines were glitched. The readbacks for Tevatron devices were gone, so it was hard to diagnose the problem. It was suspected that a power feeder had tripped off.

At 05:55, it was confirmed by the FESS electrician that Tevatron conventional power feeder 45 had tripped off.

All machines were impacted, the Linac Klystron stations, the Booster anode supplies tripped, the Tevatron and cryo, and MI devices. There are way too many things to list hear. On a positive note, the stack of PBars was not lost. (~91 mA)

The day was spent making accesses into machines and trying to recover. The Tevatron lost helium, so it will need to be cooled down. There were also vacuum problems that needed to be repaired.

By 14:06, the LINAC was up and experts were establishing beam.

By 14:45, events were put in the timeline to tuneup Booster.

The feeder fault was isolated late in the day and the power configuration for the Tevatron was restored to a mode which will allow operation. CHL and others were informed before switching in case a glitch occurred.

By 18:09, the F-Sector keys were returned and Main Injector turn on could begin. Within 10 minutes, beam was running to MiniBooNe. An access into F-Sector for a cryo readback repair had beam off to MiniBooNe until about 20:30.

At about 22:55, beam was established to the switchyard dump in preparation for sending it to the Meson experiments. Beam was to the experiments about hour later. Problems caused by the power glitch were found and repaired during this process.

Friday May 28
The midnight shift began with a stack of 102 mA and with beam operation in Recycler, Switchyard, and MiniBooNe. The Tevatron cooldown was underway.

Small problems occurred in the running machines, but nothing major.

By 06:30, cryo reported that the cooldown of the Tevatron ring was complete. Four more hours were needed for compressor pumpdowns to finish.

The Meson Center experiment dropped their interlocks for the day.

D0 and CDF will remain in access until later in the morning.

If things continue to go well, Tevatron startup may begin around noon today.

Today & Weekend Plans
The Tevatron should be cold later this morning. After work is completed in the tunnels and experimental halls, we will turn on the Tevatron for checkout. We will continue to run to MiniBooNe, SY120, and stack throughout the day. Once Tevatron checkout is complete, a store will follow. The plan is smooth operation with stacking and store over the holiday weekend.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 5/28/2004   email Fermilab